OpenStudy (anonymous):

Rewrite each function to make it easy to graph using transformations 1) \[y=\sqrt{9x-63}+4\] 2) \[y=\sqrt[3]{8x-64}-5\]

OpenStudy (anonymous):

@AnimalAin @dpaInc @dumbcow @foolformath help plz

OpenStudy (dumbcow):

1) factor out a 9 2) factor out a 8

OpenStudy (anonymous):

what do you mean?

OpenStudy (dumbcow):

factor those numbers out of the radical 9x-63 = 9(x-7) ...

OpenStudy (anonymous):

can you do an example

OpenStudy (dumbcow):

\[\sqrt{12x-20} = \sqrt{4(3x-5)}=\sqrt{4}\sqrt{3x-5} =2\sqrt{3x-5}\]